Find sin2θ if secθ=-4/3 and 90° <θ<180°

Answer by lwsshak3(11628)

Find sin2θ if secθ=-4/3 and 90° <θ<180°
secθ=1/cosθ=-4/3
cosθ=-3/4
sinθ=√(1-cos^2θ)=√(1-(9/16))=√(7/16)=√7/4
sin2θ=2sinθcosθ=2*√7/4*-3/4=-6√7/16